OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (3, 5) and has a slope of 2? A. y+3=2(x+5) B. y-3=2(x-5) C. y-5=2(x-3) D. y-5=-2(x-3)

OpenStudy (anonymous):

@animal_lover36

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@Rainbow_dashie

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

@terminally_capricous

OpenStudy (anonymous):

@Rabya!

OpenStudy (anonymous):

@Bronybro

OpenStudy (anonymous):

@Alfa_female

OpenStudy (anonymous):

any idea?

OpenStudy (rabya!):

c i think

OpenStudy (campbell_st):

well the point slope formula is \[y - y_{1} = m ( x - x_{1}) \] where m = slope and (x1, y1) is the given point you are given m = 2 and point (3, 5) just substitute and you'll have the answer

OpenStudy (anonymous):

what is substitute

OpenStudy (campbell_st):

replace the letter with a number

OpenStudy (anonymous):

ok